# WOMD-Reasoning: A Large-Scale Dataset and Benchmark for Interaction and Intention Reasoning in Driving

### v2

@@ Abstract
We propose Waymo Open Motion Dataset-Reasoning (WOMD-Reasoning), a comprehensive large-scale dataset with 3 million Q&As built on WOMD focusing on describing and reasoning interactions and intentions in driving scenarios. Existing language datasets for driving primarily capture interactions caused by close distances. However, interactions induced by traffic rules and human intentions, which can occur over long distances, are yet sufficiently covered. To address this, WOMD-Reasoning presents by far the largest multi-modal Q&A dataset on real-world driving scenarios, covering a wide range of driving topics from map descriptions and motion status descriptions to narratives and analyses of agents’ interactions, behaviors, and intentions. We further introduce Motion-LLaVA, a motion-language model fine-tuned on the proposed dataset with robust interaction reasoning capabilities. We benchmark its performance across various configurations including different input modalities, reasoning techniques, and network architectures. The robust, diverse, and multi-modal nature of WOMD-Reasoning highlights its potential to advance future autonomous driving research and enable a broad range of applications. The dataset and its vision modal extension are available at https://waymo.com/open/download, and the codes & prompts to build it are available at https://github.com/yhli123/WOMD-Reasoning.
Figure 1: Two examples of interaction descriptions and reasoning in WOMD-Reasoning dataset. (a) This example captures the traffic rule-induced interaction between the ego agent and agent #0, attributing it correctly to the stop signs. (b) This example shows the traffic light-controlled yielding interaction between the ego agent and agent #1. The front-view visualization is created by importing the traffic scenario into MetaDrive simulator [24].
@@ 1 Introduction
Language-guided driving datasets and models have drawn much attention in autonomous driving. For example, methods based on large language models (LLMs) [50, 4] improve explainability, controllability, and performance of driving tasks [7, 14, 30, 57, 23, 2, 36, 26, 9, 29, 60, 16, 11, 17, 53].
@@
Various language-based datasets [45, 37, 20, 41, 30, 18, 65, 48, 2, 36, 26, 29] have been proposed to fine-tune these LLMs. However, due to the complexity and variability of driving scenarios and driver behaviors, existing datasets often focus on describing specific scenes without reasoning about the interactions, like traffic rules and human intentions behind them, which are vital for the safety-critical decision-making in autonomous driving systems [54, 62, 63]. For example, in BDD-X [20], interactions are often limited in spatial relationships, e.g., ‘The car moves back into the left lane because the school bus in front of it is stopping.’; while in another example, DriveLM [45] attributes interaction to ‘keep a safe distance’ without further detailed explanations. As a result, most existing works are either insufficient in quantity or only cover interactions induced by proximity, as shown in our analysis in Table 1.
Considering the critical importance and difficulty of interaction modeling for autonomous driving [12, 39], it is essential to incorporate comprehensive interactive and behavioral reasoning for language-based driving. Real-world interactions are often driven by human knowledge and behavior, such as traffic rules, signals, and intentions like overtaking. However, these interactions are rarely included in previous datasets, leading to suboptimal performance of LLMs fine-tuned on these datasets. Additionally, most datasets in driving are not large enough to support multi-modal fine-tuning across diverse tasks such as scene description, prediction, and planning.
@@
This work introduces WOMD-Reasoning dataset, a large-scale multi-modal dataset centered on language Q&As based on WOMD [13], with a specific focus on interaction descriptions and reasoning. To incorporate interactions induced by traffic rules and human intentions, we build an automated data curation pipeline by prompting ChatGPT [4, 34] with a set of well-designed prompts. Together with a rule-based translator to convert motion data into language, we create an automated pipeline for generating the language dataset and achieve an approximate accuracy rate of over 90%. We then manually verify and correct a subset to provide a high-quality set suitable for supervised fine-tuning (SFT) with LLMs. To the best of our knowledge, WOMD-Reasoning is currently the largest real-world language dataset for driving. WOMD-Reasoning dataset covers interactions induced by traffic rules and human intentions with 409k Q&As, like those shown in Figure 1. As indicated in Table 1, it contains the largest number of total Q&As and interaction-specific Q&As, supporting a wide range of applications including scene description, prediction, and planning. Additionally, to facilitate multi-modal learning across motion, language, and vision modalities, for each motion data, we render both the corresponding bird’s eye view and front view images by the MetaDrive simulator [24].
Based on WOMD-Reasoning, we build our Motion-LLaVA upon LLaVA [28], which can accurately answer driving-related questions. To comprehensively absorb detailed motion information, we further replace LLaVA’s vision encoder with a motion vector encoder [51], a motion prediction model. With this design and a set of effective training and inference strategies ablated, Motion-LLaVA achieves high accuracy in both driving scenario understanding and interaction analysis, outperforming baselines by large margins. We further benchmark several settings including multi-modal model fine-tuning strategies on WOMD-Reasoning to show the advantage of Motion-LLaVA. It unleashes the analytic abilities of motion language models to provide in-depth descriptions and reasoning on driving scenarios.
@@
  • We propose WOMD-Reasoning, the largest multi-modal dataset with 3 million Q&A pairs centered on interaction and intention reasoning in driving. It provides extensive insights into critical but previously overlooked interactions induced by traffic rules and human intentions.
  • We introduce Motion-LLaVA, a motion-language model designed for WOMD-Reasoning with robust capabilities, by benchmarking task performance across various configurations, including varying input modalities (vision v.s. motion), reasoning techniques (single step v.s. chain-of-thought), and network architectures (motion encoder pre-training).
  • Fine-tuned with WOMD-Reasoning, Motion-LLaVA is capable of providing detailed and insightful descriptions and interaction reasonings on various driving scenarios. Extensive evaluations with various metrics validate its wide coverage and high quality, especially in reasoning traffic rule-induced interactions.
Table 1: Comparison between WOMD-Reasoning Dataset and Previous Real-world Language Datasets for Driving Scenarios. Compared to previous datasets, WOMD-Reasoning covers comprehensive kinds of interactions, with a significantly higher size supporting a wide range of potential applications.
@@
nuScenes-QA[37] nuScenes[5] 34k 460k 0
BDD-X[20] BDD[56] 26k 26k 26k
DriveLM[45] nuScenes[5] 5k 443k $\approx$199k
DRAMA[30] DRAMA[30] 18k 102k <18k
WOMD-Reasoning WOMD [13] 63k 2,940k 409k
@@ 2 Related Work
Application of Language Modal in Autonomous Driving. In recent years, language has been widely explored in autonomous driving, as incorporating language into implicit features could enhance the explain-ability [59, 40, 46, 6, 31]. For example, Kuo et al. [22] employ language features in the trajectory prediction model, ensuring both explainability and consistency between language and predictions. Similarly, Zhong et al. [64] use language input and LLMs to generate conditions in diffusion models, thereby controlling the generation of driving scenarios. Recently, general driving agents have been developed [57, 7, 45, 32], aiming to integrate all functions—perception, prediction, and planning—into a single language agent. However, we observe that their performance in interaction analysis is not entirely satisfying. This shortcoming is largely attributed to the lack of interaction analysis in language datasets specific to driving scenarios.
@@
Language Datasets in Autonomous Driving. Language datasets for driving scenarios have recently been developed to support related LLM-based work. Due to the controllability of simulations, several simulation-based datasets have been created [7, 45]. However, simulations often fail to capture real-world interactions. To address this, recent studies [20, 41, 30] incorporate human labeling of real-world driving scenarios to ensure the inclusion of genuine interactions. Nevertheless, obtaining high-quality human labels is labor-intensive and costly, limiting the size and coverage of these datasets. Moreover, interactions driven by traffic rules and human intentions have often been overlooked. To streamline the labeling process, some studies [10, 33, 37, 45] have generated rule-based labels, although these typically cover only basic language elements like scene descriptions, leaving interaction analysis to human labelers. To minimize human labor, we first employ manual rules and then utilize ChatGPT-4 [34] to build our dataset automatically. The resulting dataset, WOMD-Reasoning, is rich in interaction details, particularly those stemming from traffic rules and human intentions. The designed automated data-curation pipeline enables the dataset to be significantly larger. We also create a high-quality human-verified subset. Preliminary human evaluation provides positive feedback with an approximate accuracy of 90%.
@@
Fine-tuning LLM for Driving Tasks. Recent years have witnessed a surge of research on leveraging LLMs and vision-language models (VLMs) for driving applications [43, 7, 19, 58, 45, 44, 31, 47]. However, handling heterogeneous inputs such as vectorized motion representation remains challenging due to their domain gaps to most existing VLMs and LLMs. Notable prior work, such as [31], designed natural language prompt structures to accommodate numerical motion data, while another approach [7] proposed training LLMs with custom-built motion vector encoders.
@@
Our Motion-LLaVA approach is inspired by the design of [7], augmenting existing LLM architectures with tailored components. In our study, we scale the training dataset by approximately 20 times and use real-world scenarios instead of simulation-based data facilitated by WOMD-Reasoning. Furthermore, we employ a motion prediction model as a motion vector encoder, utilizing its prior knowledge to streamline the training pipeline. To enhance reasoning accuracy, we adopt a Chain-of-Thought (CoT) [55] approach, allowing the model to base its reasoning on factual grounding, mitigating potential hallucinations. With WOMD-Reasoning, our Motion-LLaVA demonstrates the practicality of applying VLMs to real-world vectorized motion data to accurately answer diverse driving-related questions correctly.
@@ 3.1 Building WOMD-Reasoning Dataset
Manually constructing autonomous driving datasets with fine-grained natural-language labels regarding the interactions among road participants would involve intense human labor, which is one of the main reasons why previous datasets have insufficient interaction analysis and a very limited overall size [42]. To reduce human labor while obtaining reasonably useful data on vehicle interactions, we propose a fully automatic data-curation pipeline to label the WOMD dataset with language: We first develop a rule-based program to interpret the motion dataset, which contains trajectories and the HD map, into language; and then build a set of prompts to utilize the reasoning abilities of ChatGPT-4 [34] to generate interaction analysis and reasoning, and organize the results into the target Q&A format. Details of each step are attached in the Appendix A.1. Codes for interpreting the motion dataset as well as the full prompts for ChatGPT-4 are available at https://github.com/yhli123/WOMD-Reasoning.
@@
Our analysis process is performed on Microsoft Azure ChatGPT-4-Turbo API, which costs around 12,750 USD. The details of WOMD-Reasoning will be provided in Part 4.
3.2 Motion-LLaVA: an Multi-modal Fine-tuning Method on WOMD-Reasoning
@@
To showcase the application of WOMD-Reasoning dataset, we perform a set of multi-modal fine-tuning experiments on it to build our Motion-LLaVA. LLaVA [28] is chosen as the pre-trained multi-modal model, in order to accommodate motion data input and text prompts input at the same time, where the vision encoder should be adapted to accommodate motion data inputs. Previous work [7] directly attended vectorized motion data with learnable queries to aggregate motion information. Inspired by [44], which trained its vision encoder with future trajectories as supervision, we utilize encoders from motion prediction models for encoding motion data. Specifically in our work, we adopted the encoder component of the motion prediction model MultiPath++ [51] as the motion vector encoder for the LLM. Ablations in Table 5 prove the helpfulness of knowledge prior in this encoder for fine-tuning. We also ablate some detailed fine-tuning strategies in Appendix A.7. Our Motion-LLaVA fine-tuning pipeline, summarized as the Motion-LLaVA method, is illustrated in Figure 3.
@@
During training, we unfreeze all components including the motion vector encoder, and train on all fields of Q&A pairs in WOMD-Reasoning simultaneously to avoid the potential catastrophic forgetting. To prevent Motion-LLaVA from forgetting its inherent LLM’s knowledge, we mix up a few questions irrelevant to the current driving scenario into the training process. [61] During inference, we adopt Chain-of-Thought (CoT) Reasoning strategy [55] to reduce the hallucination. Specifically, we let Motion-LLaVA first answer the facts questions, which include questions in map environment, ego, and other agents’ motion status. Then, we aggregate these answers and feed them into Motion-LLaVA as the contexts for answering reasoning questions, which refers to questions in interactions and intentions analysis. Examples of these contexts can be found in the Appendix A.5. We ablate in Part 5.4 that such a CoT strategy can effectively improve the quality of answering reasoning questions.
4 WOMD-Reasoning Dataset

### v3

# WOMD-Reasoning: A Large-Scale Dataset for Interaction Reasoning in Driving

@@ Abstract
Language models uncover unprecedented abilities in analyzing driving scenarios, owing to their limitless knowledge accumulated from text-based pre-training. Naturally, they should particularly excel in analyzing rule-based interactions, such as those triggered by traffic laws, which are well documented in texts. However, such interaction analysis remains underexplored due to the lack of dedicated language datasets that address it. Therefore, we propose Waymo Open Motion Dataset-Reasoning (WOMD-Reasoning), a comprehensive large-scale Q&As dataset built on WOMD focusing on describing and reasoning traffic rule-induced interactions in driving scenarios. WOMD-Reasoning also presents by far the largest multi-modal Q&A dataset, with 3 million Q&As on real-world driving scenarios, covering a wide range of driving topics from map descriptions and motion status descriptions to narratives and analyses of agents’ interactions, behaviors, and intentions. To showcase the applications of WOMD-Reasoning, we design Motion-LLaVA, a motion-language model fine-tuned on WOMD-Reasoning. Quantitative and qualitative evaluations are performed on WOMD-Reasoning dataset as well as the outputs of Motion-LLaVA, supporting the data quality and wide applications of WOMD-Reasoning, in interaction predictions, traffic rule compliance plannings, etc. The dataset and its vision modal extension are available on https://waymo.com/open/download/. The codes & prompts to build it are available on https://github.com/yhli123/WOMD-Reasoning.
Machine Learning, ICML
Figure 1: Examples of traffic rule-induced interactions in WOMD-Reasoning dataset. (a) captures the traffic rule-induced interaction between the ego agent and agent #0, attributing it correctly to the stop signs. (b) shows the traffic light-controlled yielding interaction between the ego agent and agent #1. The front-view visualization is created by MetaDrive simulator (Li et al., 2022)
@@ 1 Introduction
Language-guided models improve explainability, controllability, and performance of autonomous driving tasks (Touvron et al., 2023; Brown et al., 2020; Chen et al., 2023; Fu et al., 2023; Malla et al., 2023; Xu et al., 2024a; Li et al., 2024a; Arai et al., 2024; Park et al., 2024; Li et al., 2024b; Cui et al., 2024; Ma et al., 2024; Yuan et al., 2024; Greer & Trivedi, 2024; Ding et al., 2024a; Han et al., 2024; Wang et al., 2024b), owing to their enormous text-based knowledge acquired through the pre-training. To help language models elevate abilities on driving-related tasks, various language-based datasets (Sima et al., 2023; Qian et al., 2024; Kim et al., 2018; Sachdeva et al., 2024a; Malla et al., 2023; Inoue et al., 2024; Zhou et al., 2024; Tian et al., 2024b; Arai et al., 2024; Park et al., 2024; Li et al., 2024b; Ma et al., 2024) have been proposed to fine-tune these LLMs. However, due to the complexity and variations in summarizing traffic rule-based interactions in real driving scenarios, existing datasets often focus on interactions based on spatial proximity. For example, in BDD-X (Kim et al., 2018), interactions are often limited to the spatial blocking, like ‘The car moves back into the left lane because the school bus in front of it is stopping.’; while in another example, DriveLM (Sima et al., 2023) attributes interaction to ‘keep a safe distance’ without further detailed explanations. Likewise, most existing works only cover interactions induced by proximity, as shown in our analysis in Table 1. While reasoning about many non-proximate interactions, like those caused by traffic rules, are vital for the safety-critical decision-making in autonomous driving systems (Wang et al., 2022; Zhang et al., 2022; Zhao et al., 2024; Ding & Zhao, 2023; Roelofs et al., 2022), these interactions are rarely included in previous datasets, leading to suboptimal performance of LLMs fine-tuned on these datasets. Additionally, most datasets in driving are not large enough to support multi-modal fine-tuning across diverse tasks such as scene description, prediction, and planning.
@@
For the sake of interaction descriptions and reasoning, this work introduces WOMD-Reasoning dataset, a large-scale multi-modal dataset centered on language Q&As based on WOMD (Ettinger et al., 2021). To incorporate interactions induced by traffic rules and human intentions, we build an automated data curation pipeline by prompting ChatGPT (Brown et al., 2020; OpenAI et al., 2024) with a set of well-designed prompts. Together with a rule-based translator to convert initial motion data into language, we create an automated pipeline to generate the language dataset and achieve an approximate accuracy rate of 9̃0%. WOMD-Reasoning dataset covers interactions induced by traffic rules and human intentions with 409k Q&As, like those shown in Figure 1. As indicated in Table 1, WOMD-Reasoning contains the highest number of total Q&As and interaction-specific Q&As, supporting a wide range of applications, including scene description, prediction and planning. Figure 2 further demonstrates its uniqueness of interaction descriptions in WOMD-Reasoning with vocabulary statistics. In addition to comprehensive and in-depth interactions, WOMD-Reasoning is also the largest language dataset for real-world driving data, to the best of our knowledge and by its publication. Additionally, to facilitate multi-modal driving-related tasks like training end-to-end driving models, in addition to the LiDAR (Chen et al., 2024) info provided by Waymo, we render both the corresponding bird’s eye view and front view videos of covered driving scenarios by the MetaDrive simulator (Li et al., 2022), which are attached to the dataset.
To showcase the applications of WOMD-Reasoning, we fine-tune our Motion-LLaVA, which is built upon LLaVA (Liu et al., 2024), on WOMD-Reasoning. With careful design and a set of effective training and inference strategies ablated, Motion-LLaVA achieves high accuracy in both driving scenario understanding and interaction analysis, which unequivocally advocate the effectiveness of WOMD-Reasoning in boosting the interaction prediction abilities of driving language models.
@@
  • We propose WOMD-Reasoning, the largest multi-modal dataset with 3 million Q&A pairs centered on interaction reasoning in driving. It provides extensive insights into critical but previously overlooked interactions induced by traffic rules and human intentions. Quantitative and qualitative evaluations verify its wide coverage in interaction descriptions and reasoning, as well as its accuracy.
  • Fine-tuned with WOMD-Reasoning, our Motion-LLaVA is capable of providing detailed and insightful interaction predictions on various driving scenarios, supporting the effectiveness of WOMD-Reasoning. Extensive evaluations with various metrics validate the wide coverage and high quality of the answers of Motion-LLaVA fine-tuned on WOMD-Reasoning, ranging from driving scenario descriptions to traffic rule-compliant planning.
  • To optimize the performance of Motion-LLaVA, a motion-language model designed for WOMD-Reasoning with robust capabilities, we benchmark task performance across various configurations, including input modalities, reasoning techniques, and network architectures.
Table 1: Comparison between WOMD-Reasoning Dataset and Previous Real-world Language Datasets for Driving. WOMD-Reasoning covers more comprehensive kinds of interactions, with a significantly higher size supporting a wide range of applications.
@@
nuScenes-QA(Qian et al., 2024) nuScenes(Caesar et al., 2020) 34k 460k 0
BDD-X(Kim et al., 2018) BDD(Xu et al., 2017) 26k 26k 26k
DriveLM(Sima et al., 2023) nuScenes(Caesar et al., 2020) 5k 443k $\approx$199k
Rank2Tell(Sachdeva et al., 2024a) Rank2Tell(Sachdeva et al., 2024a) 116 >116 >116
DRAMA(Malla et al., 2023) DRAMA(Malla et al., 2023) 18k 102k <18k
NuInstruct(Ding et al., 2024b) nuScenes(Caesar et al., 2020) 850 91k <46k
nuCaption(Yang et al., 2025) nuScenes(Caesar et al., 2020) <1k 420k $\approx$140k
Tod3Cap(Jin et al., 2024) nuScenes(Caesar et al., 2020) 850 $\approx$2,300k 0
WOMD-Reasoning WOMD (Ettinger et al., 2021) 63k 2,940k 409k
@@ 2 Related Work
Application of Language Modal in Autonomous Driving. In recent years, language has been widely explored in autonomous driving, as incorporating language into implicit features could enhance the explainability (Yang et al., 2023; Roh et al., 2020; Sriram et al., 2019; Chen et al., 2019; Mao et al., 2023a; Zhang et al., 2024). For example, Kuo et al. (Kuo et al., 2022) employ language features in the trajectory prediction model, ensuring both explainability and consistency between language and predictions. Similarly, Zhong et al. (Zhong et al., 2023) use language input and LLMs to generate conditions in diffusion models, thereby controlling the generation of driving scenarios. Recently, general driving agents have been developed (Xu et al., 2024a; Chen et al., 2023; Sima et al., 2023; Mao et al., 2023b), aiming to integrate all functions—perception, prediction, and planning—into a single language agent. However, we observe that their performance in interaction analysis is not entirely satisfying, especially in non-proximal interactions like those induced by traffic rules. This shortcoming is largely attributed to the lack of interaction analysis in language datasets specific to driving scenarios.
@@
Language Datasets in Autonomous Driving. Language datasets for driving scenarios have recently been developed to support related LLM-based work. Due to the controllability of simulations, several simulation-based datasets have been created (Chen et al., 2023; Sima et al., 2023). However, simulations often fail to capture real-world interactions. To address this, recent studies (Kim et al., 2018; Sachdeva et al., 2024a; Malla et al., 2023; Wang et al., 2024a) incorporate human labeling of real-world driving scenarios to ensure the inclusion of genuine interactions. Nevertheless, obtaining high-quality human labels is labor-intensive and costly, limiting the size and coverage of these datasets. Moreover, interactions driven by traffic rules and human intentions have often been overlooked. To streamline the labeling process, some studies (Deruyttere et al., 2019; Nie et al., 2023; Qian et al., 2024; Sima et al., 2023) have generated rule-based labels, although these typically cover only basic language elements like scene descriptions, leaving interaction analysis to human labelers. To minimize human labor, we first employ manual rules and then utilize ChatGPT-4 (OpenAI et al., 2024) to build our dataset automatically. The resulting dataset, WOMD-Reasoning, is rich in interaction details, particularly those stemming from traffic rules and human intentions. The designed automated data-curation pipeline enables the dataset to be significantly larger. We also create a high-quality human-verified subset. Preliminary human evaluation provides positive feedback with an approximate accuracy of 90%.
@@
Fine-tuning LLM for Driving Tasks. Recent years have witnessed a surge of research on leveraging LLMs and vision-language models (VLMs) for driving applications (Seff et al., 2023; Chen et al., 2023; Jin et al., 2023; Xu et al., 2024b; Sima et al., 2023; Shao et al., 2023; Mao et al., 2023a; Tian et al., 2024a). However, handling heterogeneous inputs such as vectorized motion representation remains challenging due to their domain gaps to most existing VLMs and LLMs. Notable prior work, such as (Mao et al., 2023a), designed natural language prompt structures to accommodate numerical motion data, while another approach (Chen et al., 2023) proposed training LLMs with custom-built motion vector encoders.
@@
Our Motion-LLaVA approach is designed for benchmarking the performance of WOMD-Reasoning. Inspired by the design of (Chen et al., 2023), Motion-LLaVA augmenting existing LLM architectures with tailored components. We employ a motion prediction model as a motion vector encoder, utilizing its prior knowledge to streamline the training pipeline. To enhance reasoning accuracy, we adopt a Chain-of-Thought (CoT) (Wei et al., 2022) approach, allowing the model to base its reasoning on factual grounding, mitigating potential hallucinations. Based on WOMD-Reasoning, our Motion-LLaVA demonstrates the practicality of applying VLMs to real-world vectorized motion data to accurately answer diverse driving-related questions correctly.
@@ 3 Method
In this part, we first introduce how to build the WOMD-Reasoning dataset. To verify the effectiveness of WOMD-Reasoning, we introduce Motion-LLaVA, a multi-modal fine-tuning approach, to build motion language models upon WOMD-Reasoning. The Motion-LLaVA method as well as the fine-tuning process are also included here.
@@ 3.1 Building WOMD-Reasoning Dataset
Manually constructing autonomous driving datasets with fine-grained natural-language labels regarding the interactions among road participants would involve intense human labor, which is one of the main reasons why previous datasets have insufficient interaction analysis and a very limited overall size (Sachdeva et al., 2024b). To reduce human labor while obtaining reasonably useful data on vehicle interactions, we propose a fully automatic data-curation pipeline to label the WOMD dataset with language: We first develop a rule-based program to interpret the motion dataset, which contains trajectories and the HD map, into language; and then build a set of prompts to utilize the reasoning abilities of ChatGPT-4 (OpenAI et al., 2024) to generate interaction analysis and reasoning, and organize the results into the target Q&A format. Details of each step are attached in the Appendix A.1. Codes for interpreting the motion dataset as well as the full prompts for ChatGPT-4 are available in supplemental materials.
@@
Our analysis is performed on Microsoft Azure ChatGPT-4-Turbo API, which costs around 12,750 USD. The details of WOMD-Reasoning will be provided in Part 4.
3.2 Fine-tuning Multi-modal Model on WOMD-Reasoning
@@
To showcase the application of WOMD-Reasoning dataset, we use our Motion-LLaVA approach to fine-tune upon LLaVA (Liu et al., 2024). LLaVA is chosen to accommodate motion data and text prompts input at the same time, with the vision encoder adapted to a motion data encoder. Inspired by (Shao et al., 2023), we utilize encoders from motion prediction models MultiPath++ (Varadarajan et al., 2021) for encoding motion data. Ablations in Table 15 prove the usefulness of this encoder. Some detailed fine-tuning strategies are in Appendix A.10. Details on how we process agent IDs in Motion-LLaVA are listed in Appendix A.3. The overall Motion-LLaVA pipeline is illustrated in Figure 4.
@@
During training, we unfreeze all components including the motion vector encoder, and train on all Q&A pairs in WOMD-Reasoning simultaneously to avoid the potential catastrophic forgetting. To prevent Motion-LLaVA from forgetting its inherent LLM’s knowledge, we mix up a few questions irrelevant to the current driving scenario into the training process (Zhai et al., 2023). During inference, we adopt Chain-of-Thought (CoT) Reasoning strategy (Wei et al., 2022) to reduce the hallucination. Specifically, we first prompt Motion-LLaVA to answer the Factual questions, which include questions in map environment, ego, and other agents’ motion status. Then, we aggregate these answers and feed them into Motion-LLaVA as the contexts for answering Reasoning questions in interactions and intentions analysis. Examples of these contexts can be found in the Appendix A.8. We ablate in Part A.10 that such a CoT strategy can effectively improve the quality of answering reasoning questions.
Specifically, we take LLaVA-v1.5-7b (Liu et al., 2024) as the pre-trained VLM. Since MultiPath++ (Varadarajan et al., 2021) did not release their codes nor checkpoints, we take the implementations by (Konev, 2022). Our multi-modal fine-tuning takes 2 GPU days ($\approx$ 1 day on 2xNVIDIA A6000 GPUs) to train 1 epoch on the entire training set of WOMD-Reasoning. For quantitative evaluations, we randomly select $\approx$ 1,000 QA pairs from the validation set of WOMD-Reasoning.
To benchmark Motion-LLaVA model, we utilize a suite of standard language metrics, including BLEU-4 (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005), CIDEr (Vedantam et al., 2015), SPICE (Anderson et al., 2016), and ROUGE-L (Lin, 2004), along with GPT Score (Radford et al., 2018) to assess the quality of the generated content. Details on the prompts used to calculate the GPT Score can be found in Appendix A.5.
4 WOMD-Reasoning Dataset Specifications